# On minimizing cyclists’ ascent times

### v2

@@ 1 Introduction
In this article, we formulate a strategy to minimize the time of climbing an uphill with a bicycle, which is the goal for timetrial competitions composed of an uphill; in Italy, they are referred to with the specific name of a cronoscalata. We consider that the ascent profile, including distance and steepness, is known. The weight of the cyclist, together with the bicycle, as well as fitness level, quantified by the power output, are also known. With these — and other information, such as the air, rolling and drivetrain resistances — we formulate the aforementioned strategy. The formulation is based on power considerations, which — for the last couple of decades — has been the key metric in training and race preparation. Notably, most modern racing bicycles have power meters.
We begin this article by invoking a phenomenological model to estimate the power required to maintain a given speed. This model is analogous to the one used and discussed by Bos et al. (2023, 2024). Subsequently, we seek the riding strategy that results in the least ascent time, which — under stated constraints and assumptions — we show to be constant speed. We proceed to exemplify this result for several distinct ascent profiles. We conclude by discussing the obtained result, comparing it to another plausible strategy: constant power, and commenting on possible future developments. In the appendices, we provide a method for imposing an optimization constraint, namely, the value of the available average power, justify an aspect of our model simplification, namely, neglecting the power used to increase kinetic energy, examine the effect of imposing the second constraint, namely, the value of the available maximum power, and comment on numerical optimization.
@@ 3.1 Formulation
To determine the optimal strategy for minimizing the ascent time, we allow the instantaneous power to vary within the constraint of a given average power to be specified a priori based on the cyclist’s capacities for the expected duration of the effort. This power can be obtained from the cyclist’s power profile (e.g., Leo et al., 2022, Figure 2), as discussed in Appendix A. We assume the ascent to be monotonic, and model the hill by $N$ straight-line segments. Furthermore, we assume that each segment is traversed at a constant speed and we allow speed to be discontinuous between segments. [^6] The discontinuous speed results in a simplified model that is easier to understand, yet captures the essence of the physics of more sophisticated models. In particular, we have analyzed models with continuous speed — for which there does not seem to be an analytical solution — and have confirmed by numerical experimentation that the conclusions remain substantially unchanged. In the limit — as $N\rightarrow\infty$ — any realistic hill as well as any speed profile can be approximated with arbitrary accuracy. Yet, our argument presented in this section is independent of the value of $N$.
@@ 3.4 Ascent-time surface
Figure 2: Two-segment ascent and grade profile
@@
4 Application to arbitrary ascent profiles
4.1 Ascent profiles
@@
Figure 4: Ascent and grade profiles of varying concavity
As stated in Section 3, the method presented therein can be applied to any uphill by discretizing it into $N$ straight-line segments. In this section, we apply the method to various ascent profiles. To compare the optimal strategy of constant speed and variable power with another strategy, we include the ascent times under the assumption of constant power and variable speed.
To generate continuous ascent profiles, which can be sampled with arbitrary resolution, we consider $(n+1)$ nodes in $x\,G$-plane, where $x$ corresponds to the horizontal distance and $G$ is the grade, which is the tangent of the slope angle. We subdivide the distance with strictly increasing coordinates, $x_{0}<x_{1}<\cdots<x_{n}$. Along the $i$th segment, we define the grade as a piecewise linear interpolation of its endpoints,
@@ 4.2 Speed
Table 1: Ascent times for five ascent profiles
@@ 4.3 Times
Figure 5: Power and speed corresponding to the five ascent profiles in Figures (4(a)) – (4(e)). Black lines correspond to the constant-power solution and grey lines to the constant-speed solution; for profile (4(a)), both have the same solution. In contrast to Figure 4, the horizontal axis is the distance travelled, not its horizontal component.
@@ 5 Discussion and conclusions
Minimizing the ascent time is equivalent to maximizing — for a given uphill — the corresponding mean ascent velocity (VAM: velocità ascensionale media), which has typically units of vertical metres per hour. Our approach is valid, in general, given an arbitrary uphill. For a VAM problem, however, we would also look for a specific ascent profile to maximize the speed of gaining altitude, as discussed by Bos et al. (2021).
@@ Acknowledgements
We wish to acknowledge David Dalton, for his proofreading, Elena Patarini, for her graphic support, and Roberto Lauciello, for his artistic contribution. Also, we thank Stefano De Marchi for shepherding the peer-review process and the anonymous reviewer for insightful comments.
@@ Appendix B Power estimate for changes in kinetic energy
To gain an insight into the power expended to increase kinetic energy, we use the fact that — for conservative forces — we need to consider only the initial and final states. As we see in the right-hand side column of Figure 5 — and considering the cases of constant power — the same speed at the bottom and at the top of the ascent corresponds to profiles (4(a)), (4(d)), (4(e)) of Figure 4. As shown in the right-hand side column of Figure 4, the same speed — for constant power — is tantamount to the same slope. For these profiles, the change of kinetic energy is zero, since
@@
Let us consider profile (4(c)) in Figure 4, which is steeper at the bottom than at the top. In the constant-power case, the speed at the bottom is $V_{0}=2.1175\,{\rm m/s}$ and the speed at the top is $V_{N}=11.2361\,{\rm m/s}$. Hence, the change in kinetic energy is
@@
To quantify this statement, let us return to profile (4(c)) and scale the horizontal and vertical distances by $1/2$; in other words, we specify $x_{n}=500\,{\rm m}$ and $y_{n}=50\,{\rm m}$, which maintains the overall grade percentage of $\overline{\mathcal{G}}=10\%$. Also, the speeds at the bottom and top do not change. However, the ascent time is halved to $T=131.8770\,{\rm s}$ and, hence, the $P_{K}$ is doubled to $P_{K}=32.9764\,{\rm W}$. If we double the horizontal and vertical distances, specifying $x_{n}=2000\,{\rm m}$ and $y_{n}=200\,{\rm m}$, which also maintains the overall grade percentage, the speeds at the bottom and top do not change but the ascent time is doubled to $T=527.5078\,{\rm s}$ and, hence, the $P_{K}$ is halved to $P_{K}=8.2441\,{\rm W}$, which is two orders of magnitude less than the specified average power.
@@ Appendix D Numerical optimization considerations
P1. [Initialize.] Set segment speeds to the constant speed, $V_{j}=V$, that corresponds to the global minimum of the ascent time, where the only constraint is the average power, using coefficients (17) in solution (18), set $\mathcal{J}=\emptyset$, and initialize a counter $i=1$.
@@
P3. [Update constant speed.] Recalculate $V_{j}=V$ for $j\not\in\mathcal{J}$, using $p$ from expression (C.5) and $q$ from expression (17), iterate the counter $i=i+1$, and return to P2.

### v3

@@ 1 Introduction
In this article, we formulate a strategy to minimize the time of climbing an uphill with a bicycle, which is the goal for timetrial competitions composed of an uphill; in Italy, they are referred to with the specific name of a cronoscalata. We consider that the ascent, including distance and steepness, is known. The weight of the cyclist, together with the bicycle, as well as fitness level, quantified by the power output, are also known. With these — and other information, such as the air, rolling and drivetrain resistances — we formulate the aforementioned strategy. The formulation is based on power considerations, which — for the last couple of decades — has been the key metric in training and race preparation. Notably, most modern racing bicycles have power meters.
We begin this article by invoking a phenomenological model to estimate the power required to maintain a given speed. This model is analogous to the one used and discussed by Bos et al. (2023, 2024). Subsequently, we seek the riding strategy that results in the least ascent time, which — under stated constraints and assumptions — we show to be constant speed. We proceed to exemplify this result for several distinct ascents. We conclude by discussing the obtained result, comparing it to another plausible strategy: constant power, and commenting on possible future developments. In the appendices, we provide a method for imposing an optimization constraint, namely, the value of the available average power, justify an aspect of our model simplification, namely, neglecting the power used to increase kinetic energy, examine the effect of imposing the second constraint, namely, the value of the available maximum power, comment on numerical optimization, and discuss an ad hoc constant-ascent-speed strategy.
@@ 3.1 Formulation
To determine the optimal strategy for minimizing the ascent time, we allow the instantaneous power to vary within the constraint of a given average power to be specified a priori based on the cyclist’s capacities for the expected duration of the effort. This power can be obtained from the cyclist’s power profile (e.g., Leo et al., 2022, Figure 2), as discussed in Appendix A. We assume the ascent to be monotonic, and model the hill by $N$ straight-line segments. Furthermore, we assume that each segment is traversed at a constant speed and we allow speed to be discontinuous between segments. [^6] The discontinuous speed results in a simplified model that is easier to understand, yet captures the essence of the physics of more sophisticated models. In particular, we have analyzed models with continuous speed — for which there does not seem to be an analytical solution — and have confirmed by numerical experimentation that the conclusions remain substantially unchanged. In the limit — as $N\rightarrow\infty$ — any realistic hill as well as any speed can be approximated with arbitrary accuracy. Yet, our argument presented in this section is independent of the value of $N$.
@@ 3.4 Ascent-time surface
Figure 2: Two-segment ascent and grade
@@
4 Application to arbitrary ascents
4.1 Ascent
@@
Figure 4: Ascent and grades of varying concavity
As stated in Section 3, the method presented therein can be applied to any uphill by discretizing it into $N$ straight-line segments. In this section, we apply the method to various ascents. To compare the optimal strategy of constant speed and variable power with another strategy, we include the ascent times under the assumption of constant power and variable speed.
To generate continuous ascents, which can be sampled with arbitrary resolution, we consider $(n+1)$ nodes in $x\,G$-plane, where $x$ corresponds to the horizontal distance and $G$ is the grade, which is the tangent of the slope angle. We subdivide the distance with strictly increasing coordinates, $x_{0}<x_{1}<\cdots<x_{n}$. Along the $i$th segment, we define the grade as a piecewise linear interpolation of its endpoints,
@@ 4.2 Speed
Table 1: Ascent times for five ascents
@@ 4.3 Time
Figure 5: Power and speed corresponding to the five ascents in Figures (4(a)) – (4(e)). Black lines correspond to the constant-power solution and grey lines to the constant-speed solution; for ascent (4(a)), both have the same solution. In contrast to Figure 4, the horizontal axis is the distance travelled, not its horizontal component.
@@ 5 Discussion and conclusions
Minimizing the ascent time is equivalent to maximizing — for a given uphill — the corresponding mean ascent velocity (VAM: velocità ascensionale media), which has typically units of vertical metres per hour. Our approach is valid, in general, given an arbitrary uphill. For a VAM problem, however, we would also look for a specific ascent to maximize the speed of gaining altitude, as discussed by Bos et al. (2021).
@@
In Appendix E, we consider an ad hoc strategy, which, in contrast to constant $V$, does not emerge from a solution of an optimization problem; its inclusion is motivated by questions from our readers. That strategy requires a constant ascent speed, which is the vertical component of $V$. As we point out in the appendix, that strategy breaks down for ascents with horizontal portions. Furthermore — even for ascents without such portions — it is much inferior to either a constant-speed or constant-power strategy.
@@ Acknowledgements
We wish to acknowledge David Dalton, for his proofreading, Elena Patarini, for her graphic support, and Roberto Lauciello, for his artistic contribution.
@@ Appendix B Power estimate for changes in kinetic energy
To gain an insight into the power expended to increase kinetic energy, we use the fact that — for conservative forces — we need to consider only the initial and final states. As we see in the right-hand side column of Figure 5 — and considering the cases of constant power — the same speed is at the bottom and at the top of ascents (4(a)), (4(d)), (4(e)) of Figure 4. As shown in the right-hand side column of Figure 4, the same speed — for constant power — is tantamount to the same slope. For these ascents, the change of kinetic energy is zero, since
@@
Let us consider ascent (4(c)) in Figure 4, which is steeper at the bottom than at the top. In the constant-power case, the speed at the bottom is $V_{0}=2.1175\,{\rm m/s}$ and the speed at the top is $V_{N}=11.2361\,{\rm m/s}$. Hence, the change in kinetic energy is
@@
To quantify this statement, let us return to ascent (4(c)) and scale the horizontal and vertical distances by $1/2$; in other words, we specify $x_{n}=500\,{\rm m}$ and $y_{n}=50\,{\rm m}$, which maintains the overall grade percentage of $\overline{\mathcal{G}}=10\%$. Also, the speeds at the bottom and top do not change. However, the ascent time is halved to $T=131.8770\,{\rm s}$ and, hence, the $P_{K}$ is doubled to $P_{K}=32.9764\,{\rm W}$. If we double the horizontal and vertical distances, specifying $x_{n}=2000\,{\rm m}$ and $y_{n}=200\,{\rm m}$, which also maintains the overall grade percentage, the speeds at the bottom and top do not change but the ascent time is doubled to $T=527.5078\,{\rm s}$ and, hence, the $P_{K}$ is halved to $P_{K}=8.2441\,{\rm W}$, which is two orders of magnitude less than the specified average power.
@@ Appendix D Numerical optimization considerations
P1. [Initialize.] Set segment speeds to the constant speed, $V_{j}=V$, that corresponds to the global minimum of the ascent time, where the only constraint is the average power, using coefficients (E.1) in solution (18), set $\mathcal{J}=\emptyset$, and initialize a counter $i=1$.
@@
P3. [Update constant speed.] Recalculate $V_{j}=V$ for $j\not\in\mathcal{J}$, using $p$ from expression (C.5) and $q$ from expression (E.1), iterate the counter $i=i+1$, and return to P2.
@@
Appendix E Constant ascent speed
Let us discuss yet another strategy, namely, a constant ascent speed, for which the vertical component of the ground speed, $V_{j}\sin\theta_{j}=:v$, is constant, for $j=1,\dots,N$. Together with ground speed and power, the ascent speed is provided by a cyclocomputer attached to the handlebars, and hence is a common performance metric. Following the procedure presented in Section 4.2, we rewrite the power constraint as $0=v^{3}+p\,v+q$, which is analogous to equation (16), with
$p=\frac{1}{\beta}\dfrac{\sum\limits_{j=1}^{N}\alpha_{j}L_{j}}{\sum\limits_{j=1}^{N}L_{j}\csc^{2}\theta_{j}}\quad\text{and}\quad q=-\frac{P_{0}}{\beta}\dfrac{\sum\limits_{j=1}^{N}L_{j}\sin\theta_{j}}{\sum\limits_{j=1}^{N}L_{j}\csc^{2}\theta_{j}}\,.$ (E.1)
However, these coefficients present a problem if $\theta_{j}=0$ and, hence, $\csc^{2}\theta_{j}\to\infty$; coefficients (E.1) tend to zero. In such cases, the problem of a constant ascent speed is ill-posed. The cubic equation becomes $v^{3}=0$ ; hence, $v=0$. If there is a point where $\theta=0$ — such as the initial or final point in ascents (4(b)) and (4(c)), respectively, or the inflection point, in ascents (4(d)) and (4(e)) — we have an asymptotic ill-posedeness. If there is a horizontal segment, we have a direct ill-posedeness.
The question of ascent time could be examined further. The series for $T$ includes terms of the form $v/\sin\theta_{j}$, which lead to indeterminate forms of $/0$ 0. However, we choose not to engage in elucidating this ill-posed problem.
Be that as it may — even if ill-posedeness can be avoided by excluding points or segments where $\theta=0$ — a constant ascent speed is not a viable strategy. For ascent (4(a)), since the grade is constant, the ground speed, $V=3.9439\,\rm m/s$, power, $P=300\,\rm W$, and ascent speed, $v=0.3924\,\rm m/s$, are constant, and result in the same ascent time, $T=254.8206\,\rm s$ , as expected in view of Table 1 and $v=V\,\sin(5.7106^{\circ})$ In general, however, the instantaneous power and ground speed corresponding to a constant ascent speed are excessively large along less steep portions of the climb in order to keep the vertical component of speed constant. Notably, this supports the choice of a consistently steep slope to maximize VAM.